\begin{thebibliography}{10}

\bibitem{AGMSV:2017}
Hannah Altmann, Elo\'isa Grifo, Jonathan Monta\~no, William T. Sanders, and Thanh Vu,
\newblock Lower bounds on projective levels of complexes,
\newblock {\em J. Algebra} {\bf 491} (2017), 343--356.
\newblock \MR{3699100}.

\bibitem{A:2018}
Yves Andr\'e,
\newblock La conjecture du facteur direct,
\newblock {\em Publ. Math. Inst. Hautes \'Etudes Sci.} {\bf 127} (2018), 71--93.
\newblock \MR{3814651}.

\bibitem{A:2010}
Luchezar L. Avramov,
\newblock Infinite free resolutions,
\newblock In {\em Six lectures on commutative algebra}, volume 331 of {\em Mod. Birkh\"auser Class.}, pages 1--118,
Birkh\"auser Verlag, Basel, 2010.
\newblock \MR{2641236}.

\bibitem{ABIM:2010}
Luchezar L. Avramov, Ragnar-Olaf Buchweitz, Srikanth B. Iyengar, and Claudia Miller,
\newblock Homology of perfect complexes,
\newblock {\em Adv. Math.} {\bf 223} (2010), no. 5, 1731--1781.
\newblock \MR{2592508}.

\bibitem{AI:2007}
Luchezar L. Avramov and Srikanth B. Iyengar,
\newblock Constructing modules with prescribed cohomological support,
\newblock {\em Illinois J. Math.} {\bf 51} (2007), no. 1, 1--20.
\newblock \MR{2346182}.

\bibitem{AIN:2018}
Luchezar L. Avramov, Srikanth B. Iyengar, and Amnon Neeman,
\newblock Big {C}ohen-{M}acaulay modules, morphisms of perfect complexes, and intersection theorems in local algebra,
\newblock {\em Doc. Math.} {\bf 23} (2018), 1601--1619.
\newblock \MR{3890961}.

\bibitem{B:2021}
Bhargav Bhatt,
\newblock Cohen-Macaulayness of absolute integral closures,
\newblock preprint, 2021, \url{https://arxiv.org/abs/2008.08070}.

\bibitem{B:2023}
Sylvain Brochard,
\newblock Independent sequences and freeness criteria,
\newblock {\em J. Algebra} {\bf 628} (2023), 486--508.
\newblock \MR{4575651}.

\bibitem{CF:2022}
Lars Winther Christensen and Luigi Ferraro,  
\newblock The Improved New Intersection Theorem Revisited,  
\newblock {\em Michigan Math. J.} {\bf 75} (2025), no. 2.  
\newblock \MR{4905032}

\bibitem{CFH:2024}
Lars Winther Christensen, Hans-Bj\o rn Foxby, and Henrik Holm,
\newblock {\em Derived Category Methods in Commutative Algebra},
\newblock Springer Monographs in Mathematics, Springer, 2024.

\bibitem{DG:2002}
W. G. Dwyer and J. P. C. Greenlees,
\newblock Complete modules and torsion modules,
\newblock {\em Amer. J. Math.} {\bf 124} (2002), no. 1, 199--220.
\newblock \MR{1879003}.

\bibitem{EG:1981}
E. Graham Evans and Phillip Griffith,
\newblock The syzygy problem,
\newblock {\em Ann. of Math. (2)} {\bf 114} (1981), no. 2, 323--333.
\newblock \MR{632842}.

\bibitem{FI:2003}
Hans-Bj\o rn Foxby and Srikanth Iyengar,
\newblock Depth and amplitude for unbounded complexes,
\newblock In {\em Commutative algebra (Grenoble/Lyon, 2001)}, volume 331 of {\em Contemp. Math.}, pages 119--137,
Amer. Math. Soc., Providence, RI, 2003.
\newblock \MR{2013162}.

\bibitem{GM:1992}
J. P. C. Greenlees and J. P. May,
\newblock Derived functors of {$I$}-adic completion and local homology,
\newblock {\em J. Algebra} {\bf 149} (1992), no. 2, 438--453.
\newblock \MR{1172439}.

\bibitem{Ha:2001}
Douglas Hanes,
\newblock Length approximations for independently generated ideals,
\newblock {\em J. Algebra} {\bf 237} (2001), no. 2, 708--718.
\newblock \MR{1816712}.

\bibitem{H:1975}
Melvin Hochster,
\newblock Big {C}ohen-{M}acaulay modules and algebras and embeddability in rings of {W}itt vectors,
\newblock In {\em Conference on commutative algebra---1975 (Queen's Univ., Kingston, ON, 1975)}, pages 106--195,
Queen's Univ., Kingston, ON, 1975.
\newblock \MR{396544}.

\bibitem{H:1983}
Melvin Hochster,
\newblock Canonical elements in local cohomology modules and the direct summand conjecture,
\newblock {\em J. Algebra} {\bf 84} (1983), no. 2, 503--553.
\newblock \MR{723406}.

\bibitem{H:2002}
Melvin Hochster,
\newblock Big {C}ohen-{M}acaulay algebras in dimension three via {H}eitmann's theorem,
\newblock {\em J. Algebra} {\bf 254} (2002), no. 2, 395--408.
\newblock \MR{1933876}.

\bibitem{HH:1991}
Melvin Hochster and Craig Huneke,
\newblock Absolute integral closures are big {C}ohen-{M}acaulay algebras in characteristic {$P$},
\newblock {\em Bull. Amer. Math. Soc. (N.S.)} {\bf 24} (1991), no. 1, 137--143.
\newblock \MR{1056558}.

\bibitem{HH:1992}
Melvin Hochster and Craig Huneke,
\newblock Infinite integral extensions and big {C}ohen-{M}acaulay algebras,
\newblock {\em Ann. of Math. (2)} {\bf 135} (1992), no. 1, 53--89.
\newblock \MR{1147957}.

\bibitem{H:1987}
Michael J. Hopkins,
\newblock Global methods in homotopy theory,
\newblock In {\em Homotopy theory (Durham, 1985)}, volume 117 of {\em London Math. Soc. Lecture Note Ser.}, pages 73--96,
Cambridge Univ. Press, Cambridge, 1987.
\newblock \MR{932260}.

\bibitem{I:1999}
S. Iyengar,
\newblock Depth for complexes, and intersection theorems,
\newblock {\em Math. Z.} {\bf 230} (1999), no. 3, 545--567.
\newblock \MR{1680036}.

\bibitem{I:2001}
Srikanth Iyengar,
\newblock Free summands of conormal modules and central elements in homotopy {L}ie algebras of local rings,
\newblock {\em Proc. Amer. Math. Soc.} {\bf 129} (2001), no. 6, 1563--1572.
\newblock \MR{1707520}.

\bibitem{IMSW:2021}
Srikanth B. Iyengar, Linquan Ma, Karl Schwede, and Mark E. Walker,
\newblock Maximal {C}ohen-{M}acaulay complexes and their uses: a partial survey,
\newblock In {\em Commutative algebra}, pages 475--500,
\newblock Springer, Cham, [2021].
\newblock \MR{4394418}

\bibitem{K:1988}
Jee Koh,
\newblock Super height of an ideal in a {N}oetherian ring,
\newblock {\em J. Algebra} {\bf 116} (1988), no. 1, 1--6.
\newblock \MR{944148}

\bibitem{L:1964}
Christer Lech,
\newblock Inequalities related to certain couples of local rings,
\newblock {\em Acta Math.} {\bf 112} (1964), 69--89.
\newblock \MR{161876}

\bibitem{L:2021}
Janina C. Letz,
\newblock Local to global principles for generation time over commutative noetherian rings,
\newblock {\em Homology Homotopy Appl.} {\bf 23} (2021), no. 2, 165--182.
\newblock \MR{4259573}

\bibitem{N:1992}
Amnon Neeman,
\newblock The chromatic tower for {$D(R)$},
\newblock With an appendix by Marcel B\"okstedt,
\newblock {\em Topology} {\bf 31} (1992), no. 3, 519--532.
\newblock \MR{1174255}

\bibitem{PS:1974}
Christian Peskine and Lucien Szpiro,
\newblock Syzygies et multiplicit\'es,
\newblock {\em C. R. Acad. Sci. Paris S\'er. A} {\bf 278} (1974), 1421--1424.
\newblock \MR{349659}

\bibitem{R:1987}
Paul Roberts,
\newblock Le th\'eor\`eme d'intersection,
\newblock {\em C. R. Acad. Sci. Paris S\'er. I Math.} {\bf 304} (1987), no. 7, 177--180.
\newblock \MR{880574}

\bibitem{V:1967}
Wolmer V. Vasconcelos,
\newblock Ideals generated by {$R$}-sequences,
\newblock {\em J. Algebra} {\bf 6} (1967), 309--316.
\newblock \MR{213345}

\end{thebibliography}

\documentclass{amsart}
\usepackage[utf8]{inputenc}
\usepackage{amssymb,bm,amsfonts, stmaryrd, amsmath, amsthm, enumerate, mathtools,shuffle,colonequals,bbm}
\usepackage[shortlabels]{enumitem}
\usepackage{hyperref}
\usepackage{xifthen}
\usepackage{diagbox}
\usepackage{caption,subcaption}
\usepackage{stackengine}
\stackMath
\usepackage[mathlines]{lineno}
\hypersetup{
 colorlinks=true,
 linkcolor=blue,
 filecolor=magenta, 
 urlcolor=black,
}
\usepackage{tikz-cd}
\tikzcdset{nodes in empty cells}

\usepackage[capitalise]{cleveref}
\crefformat{equation}{(#2#1#3)}
\crefrangeformat{equation}{(#3#1#4--#5#2#6)}
\crefmultiformat{equation}{(#2#1#3}{ \& #2#1#3)}{, #2#1#3}{ \& #2#1#3)}

\crefformat{enumi}{(#2#1#3)}
\crefrangeformat{enumi}{#3#1#4--#5#2#6}
\crefmultiformat{enumi}{#2#1#3}{ and #2#1#3}{, #2#1#3}{ and #2#1#3}

\crefname{construction}{Construction}{Constructions}
\crefname{discussion}{}{}

\newtheorem{theorem}[subsection]{Theorem}

\newtheorem{proposition}[subsection]{Proposition}
\newtheorem{corollary}[subsection]{Corollary}

\newcounter{intro}

\theoremstyle{definition}
\newtheorem{definition}[subsection]{Definition}

\theoremstyle{remark}
\newtheorem{remark}[subsection]{Remark}

\numberwithin{equation}{subsection}

\theoremstyle{plain}
\newtheorem{innercustomthm}{Theorem}
\newenvironment{customthm}[1]
  {\renewcommand{\theinnercustomthm}{#1}%
   \begin{innercustomthm}}
  {\end{innercustomthm}}

\newtheorem{innercustomprop}{Proposition}

\DeclareMathOperator{\depth}{depth}
\DeclareMathOperator{\edim}{edim}

\DeclareMathOperator{\projdim}{projdim}

\DeclareMathOperator{\Ass}{Ass}
\newcommand{\cat}[1]{{\mathsf{#1}}}

\newcommand{\Ext}[4][]{\operatorname{Ext}^{#1}_{#2}(#3,#4)}
\DeclareMathOperator{\height}{height}
\DeclareMathOperator{\bigheight}{bigheight}
\DeclareMathOperator{\superheight}{superheight}

\newcommand{\Hom}[4][]{\operatorname{Hom}^{#1}_{#2}(#3,#4)}

\DeclareMathOperator{\level}{level}

\DeclareMathOperator{\f_rank}{f-rank}

\newcommand{\set}[2]{\left\{#1 \,\middle|\, #2\right\}}

\newcommand{\susp}{\Sigma}

\DeclareMathOperator{\Spec}{Spec}

\newcommand{\Tor}[4][]{\operatorname{Tor}_{#1}^{#2}(#3,#4)}

\newcommand{\fm}{\mathfrak{m}}
\newcommand{\fn}{\mathfrak{n}}
\newcommand{\fp}{\mathfrak{p}}

\begin{document}

\title[Level Inequalities]{A lower bound on levels with \\ applications to Koszul Complexes}

\author[A. Kekkou]{Antonia Kekkou}
\address{Department of Mathematics,
University of Utah, Salt Lake City, UT 84112, U.S.A.}
\email{kekkou@math.utah.edu}

\date{\today}
\keywords{Improved New Intersection Theorem, level, Koszul complex, free rank}

\begin{abstract}
In this paper, we establish a lower bound on the level of a perfect complex with $I$-power torsion homology on positive degrees and an $I$-power torsion minimal generator for $H_0(F)$. Examples are provided to demonstrate that the bound is optimal. This result is applied to improve existing lower bounds on the level of a Koszul complex on various classes of sequences.
\end{abstract}

\maketitle
\setcounter{tocdepth}{1}

\section{Introduction}
This paper concerns certain homological invariants of finite free complexes over commutative rings. We prove the following:
\begin{customthm}{\ref{theorem_level_ineq}} Let $R$ be a commutative noetherian local ring, $I$ an ideal in $R$ and 
\begin{equation*}
F \colon 0 \longrightarrow F_n \longrightarrow F_{n-1} \longrightarrow \cdots \longrightarrow F_1 \longrightarrow F_0 \longrightarrow 0
\end{equation*}
a finite free $R$-complex with $H_0(F) \neq 0$. If $H_i(F)$ is $I$-power torsion for $i\geq 1$ and a minimal generator of $H_0(F)$ is $I$-power torsion, then the following inequality holds:
\begin{equation*}
\level^R F  \geq \dim R-\dim R/I+1.
\end{equation*}
\end{customthm}

An element is said to be \emph{$I$-power torsion} if there exists an $s>0$ such that $I^s$ annihilates it. Accordingly, a module is said to be \emph{$I$-power torsion} if each of its elements is $I$-power torsion. The $R$-level of a finite free $R$-complex measures the minimal number of mapping cones required to construct the complex from finite free modules; see \cite[Section 2]{ABIM:2010}. The $R$-level, in a way, serves as a measure of the complexity of a complex. It is bounded above by the length of the complex; see \cite[Lemma 2.5.2]{ABIM:2010}. From this observation, we deduce that \cref{theorem_level_ineq} refines \cite[Theorem 2.2]{CF:2022} due to Christensen and Ferraro, which establishes that a complex $F$ as above has length at least $\dim R-\dim R/I$.

Another closely related result is due to Avramov, Iyengar, and Neeman in \cite[Theorem 4.2]{AIN:2018}. Their result is that the $R$-level of $F$ must be at least $\height I+1$. \cref{theorem_level_ineq} extends this result, as the inequality $\dim R-\dim R/I \geq \height I$ always holds. 

These results fall under the general framework of Evans and Griffith's version of the New Intersection Theorem \cite{EG:1981}, stated by Hochster in \cite{H:1983}. Evans and Griffith's version is a generalization of the New Intersection Theorem due to Peskine and Szpiro in \cite{PS:1974}, and Roberts in \cite{R:1987}. The proof of the \cref{theorem_level_ineq} is derived by combining the proofs of two earlier versions of the New Intersection Theorem, namely those in \cite{AGMSV:2017} and \cite{CF:2022}.

\section{Preliminaries}
Let $R$ be a commutative noetherian ring. By an $R$-complex, we mean a chain complex of $R$-modules; we use lower indexing. We write $\cat{D}(R)$ for the derived category of $R$-modules, which we view as a triangulated category with the usual suspension $\susp$ acting as the translation functor. The homological supremum and infimum of an $R$-complex $M$, are denoted by
\begin{align*}
\sup H_*(M)=\sup \set{i\in \mathbb{Z}}{H_i(M)\neq 0} \\
\inf H_*(M)=\inf \set{i\in \mathbb{Z}}{H_i(M)\neq 0}.
\end{align*}
We write $K(\underline{x};M)$ for the Koszul complex on a sequence $\underline{x}=x_1, \ldots, x_n$ over an $R$-complex $M$ and $H(\underline{x}; M)$ for its homology. 

\subsection{Local Cohomology} Let $I$ be an ideal and $M$ an $R$-complex. The $I$-power torsion subcomplex of $M$ in degree $i\in \mathbb{Z}$ is defined by
\begin{equation*}
(\Gamma_I M)_i \coloneqq \set{m \in M_i}{ I^n m = 0, \text{ for some } n \geq 0}.
\end{equation*}
The corresponding right derived functor is denoted by $R\Gamma_I (M)$. The local cohomology modules of $M$ supported on $I$ are computed by
\begin{equation*}
H^i_I (M) \coloneqq H^i(R\Gamma_I (M)) \text{ for } i \in \mathbb{Z}.
\end{equation*}

\subsection{Depth} The \emph{$I$-depth} of $M$ is given by
\begin{equation*}
\depth_R(I,M) \coloneqq \inf \set{i}{H^i_I (M) \neq 0}
\end{equation*}
and it is infinity if $H^i_I (M) = 0$ for all $i$; see \cite{IMSW:2021}. In a local ring $(R, \fm)$, the $\depth_R(M)$ refers to the $\depth_R(\fm,M)$. From \cite{I:1999}, depth can also be computed using Ext groups and Koszul homology. For Ext groups, we have
\begin{equation*}
\depth_R(I,M) = \inf\set{i}{\Ext [i] {R}  {R/I} {M}  \neq 0}.    
\end{equation*} 
If a sequence $\underline{x} \coloneqq x_1, \ldots, x_n$ generates $I$, then we also have
\begin{equation} \label{depth_koszul}
\depth_R(I,M) = n - \sup\set{i}{H_i(\underline{x};M) \neq 0}.
\end{equation} 

\subsection{Auslander-Buchsbaum equality} \label{Auslander_Buchsbaum} Let $R$ be a commutative noetherian local ring. An $R$-complex $F$ is said to be \emph{perfect} if it is quasi-isomorphic to a finite free $R$-complex. For such an $F$ and any $R$-complex $M$, one has
\begin{equation*}
\depth_R(M\otimes_R^L F)=\depth_R M-\projdim_R F.    
\end{equation*}
See \cite[Theorem 2.4]{FI:2003}.

\subsection{Derived complete complexes}  \label{derived_I_complete_tensor_perfect_complex} Let $I$ be an ideal in $R$. We denote the left derived $I$-completion functor by $L\Lambda^I$. An $R$-complex $M$ is called \emph{derived $I$-complete} if the natural map
\begin{equation*}
M \overset{}{\longrightarrow} L\Lambda^I (M)    
\end{equation*}
is a quasi-isomorphism; see \cite{CFH:2024,DG:2002, GM:1992, IMSW:2021} for details.

When $F$ is a perfect $R$-complex and $M$ is a derived $I$-complete $R$-complex, then $F\otimes_R M$ is also a derived $I$-complete $R$-complex. This holds due to the following canonical map from \cite[1.10]{FI:2003}
\begin{equation*}
N \otimes_R^L L\Lambda^I M \longrightarrow L\Lambda^I (N \otimes_R^L M)\,,    
\end{equation*}
which becomes isomorphism when $N$ is a perfect $R$-complex.

\begin{proposition}
\cite[Remark 1.7]{IMSW:2021}  Let $I$ be an ideal and $M$ an $R$-complex. If $\sup H_*(M)<\infty $, then the following inequality holds
\begin{equation}\label{depth_sup_ineq}
\depth_R (I, M) \geq -\sup H_*(M)\,,
\end{equation} 
with equality if and only if $\Gamma_I (H_{s} (M))\neq 0$, where $s\coloneqq \sup H_*( M)$. 
\end{proposition}

The following result derives from the proof of \cite[Theorem 2.7]{IMSW:2021}.

\begin{proposition} \cite[Theorem 2.7]{IMSW:2021}
Let $(R, \fm)$ be a commutative noetherian local ring, $I$ an ideal of $R$ and $M$ a derived $\fm$-complete $R$-complex. Then, the following holds
\begin{equation} \label{depth_ideal_ineq}
\depth_R M \leq \depth_R (I,M) + \dim R/I .
\end{equation} 
More specifically, for every prime ideal $\fp$, the following inequality holds
\begin{equation}
\label{depth_prime_ineq}
\depth_R M  \leq \depth_{R_{\fp}} M_{\fp} + \dim R/\fp.
\end{equation}   
\end{proposition}

\subsection{Level} \label{Level} We can define the level of an $R$-complex with respect to any complex, but our focus is on the level with respect to $R$. 

\begin{definition} \cite[2.3]{ABIM:2010}
Let $M$ be an $R$-complex. The level of $M$ with respect to $R$, or just $R$-level is defined as follows
\begin{equation*}
\level^R M \coloneqq \inf
\set{ n\geq 0\ }{ \begin{gathered}
\text{ there is an exact triangle} \\
K \to L \oplus M \to N\to \susp K \\
\text{with} \level^R K=1 \text{ and } \level^R N= n-1
\end{gathered}}
\end{equation*}
where $\level^R M=0$ if $M$ is quasi-isomorphic to zero, and $\level^R M=1$ if $M$ is built out of $R$ using (de)suspensions, retracts and finite coproducts.
\end{definition}

An $R$-complex $F$ is perfect if and only if it has finite level with respect to $R$. The length of a perfect complex provides an upper bound for its $R$-level \cite[Lemma 2.5.2]{ABIM:2010}, but it can be arbitrarily larger than the $R$-level. For instance, over a regular local ring, the $R$-level of any perfect $R$-complex cannot exceed the ring's dimension, while there are perfect $R$-complexes with arbitrarily large length; see \cite[Example 5.3]{ABIM:2010}. 

The following result is a special case of \cite[Lemma 2.4 (6)]{ABIM:2010} and provides a comparison of the level after base change.

\begin{proposition} \cite[Lemma 2.4 (6)]{ABIM:2010} 
If $S$ is an $R$-algebra, given the exact functor $-\otimes^L_R S \colon \cat{D}(R) \to \cat{D}(S)$, then for any $M\in \cat{D}(R)$, the following inequality holds
\begin{equation}  \label{level_change_basis}
\level^R M\geq \level^S (M \otimes_R^L S).
\end{equation}
\end{proposition}

For an $R$-complex $M$ with non-zero homology, let $P$ be a projective resolution of $M$. For $n \in \mathbb{Z}$, we denote the nth syzygy of $M$ by $\Omega^R_n(M)\coloneqq \susp^{-n} (P_{\geq n})$. It has been shown in  \cite[Lemma 1.2]{AI:2007} that whether $H_0(\Omega^R_n(M))$ is projective is independent of the choice of the projective resolution $P$. The following proposition from \cite{AGMSV:2017} is key to the proof of the level inequality \ref{theorem_level_ineq}.

\begin{proposition} \cite[Theorem 2.1 and Remark 2.5]{AGMSV:2017} Let $M$ be an $R$-complex, with $H_i(M) =0$ for all $a <i<b$, $a, b \in \mathbb{Z}$ and $H_0(\Omega_{b-1}^R(M))$ is not projective, then
\begin{equation} \label{level_bound_ineq}
\level^R M \geq b - a + 1.
\end{equation}   
\end{proposition} 

In the preceding result, $R$ need not be noetherian. 

\subsection{Balanced big Cohen-Macaulay algebras} The proof of Theorem \ref{theorem_level_ineq} uses the existence of balanced big Cohen–Macaulay algebras. An $R$-algebra $S$ is called \emph{balanced big Cohen–Macaulay algebra} if every system of parameters for $R$, forms an $S$-regular sequence. The existence of such algebras has been proved by Hochster and Huneke in  \cite{H:1975}, \cite{H:2002}, \cite{HH:1991} and \cite{HH:1992} when $R$ is equicharacteristic or $\dim R \leq 3$, and in general by André, in his recent work on the Direct Summand Conjecture \cite{A:2018}. See also \cite{B:2021} for a different proof by Bhatt when $R$ has mixed characteristic. 

\subsection{Acknowledgments} I would like to thank my advisor, Srikanth Iyengar. I am deeply grateful to him for suggesting this problem and for his invaluable guidance, insightful discussions, detailed comments on the manuscript, and constant support throughout the development of this work. I would also like to thank Janina Letz, Des Martin, Claudia Miller and Josh Pollitz for useful conversations on the material. This work is partly supported by National Science Foundation grant DMS-200985. 

\section{The level inequalities}
Here is the Theorem from the introduction.

\begin{theorem}
\label{theorem_level_ineq}  Let $R$ be a commutative noetherian local ring, $I$ an ideal in $R$ and 
\begin{equation*}
F \colon 0 \longrightarrow F_n \longrightarrow F_{n-1} \longrightarrow \cdots \longrightarrow F_1 \longrightarrow F_0 \longrightarrow 0
\end{equation*}
a finite free $R$-complex with $H_0(F) \neq 0$. If $H_i(F)$ is $I$-power torsion for $i\geq 1$ and a minimal generator of $H_0(F)$ is $I$-power torsion, then the following inequality holds:
\begin{equation*}
\level^R F  \geq \dim R-\dim R/I+1.
\end{equation*}
\end{theorem}

\begin{proof}
From $H_0(F)\neq 0$, we have that $\level^R F \geq 1$, so it suffices to prove the inequality for when $\dim R-\dim R/I \geq 1$. Also, by replacing $F$ with its minimal free resolution, we can assume that $F$ is minimal and that $F_n\neq 0$. Next, we take a balanced big Cohen-Macaulay $R$-algebra, and we complete it with respect to $\fm$, obtaining an $\fm$-complete big Cohen-Macaulay $R$-algebra, which we denote by $S$. Set $s\coloneqq \sup H_* (F \otimes_R S)$, and take $\fp \in \Ass H_s(F \otimes_R S)$. Observe that $H(F)_{\fp}\neq 0$, which will be used later. 

We claim that the following inequality always holds
\begin{equation}
\label{inequality_with_I}
n\geq \dim R-\dim R/I +s.
\end{equation}
First, consider the case when $s=0$. It follows from $H_0(F)$ being finitely generated, Nakayama’s Lemma, and \cite[Lemma 2.2]{IMSW:2021} that each minimal generator of $H_0(F)$ gives rise to a nonzero element in $H_0(F \otimes_R S)$. By hypothesis, there exists an $I$-power torsion minimal generator of $H_0(F)$ and we can lift it to a non-zero $I$-power torsion element of $H_0(F \otimes_R S)$, meaning that  $\Gamma_I(H_0(F \otimes_R S)) \neq  0$. Therefore, by \eqref{depth_sup_ineq}, we have
\begin{equation*}
\depth_R(I, F \otimes_R S) = -\sup H_* (F \otimes_R S) = 0.    
\end{equation*}
Due to \eqref{derived_I_complete_tensor_perfect_complex}, the $R$-complex $F \otimes_R S$ is a non-zero derived $\fm$-complete $R$-complex and applying \eqref{depth_ideal_ineq} yields
\begin{equation*}
\depth_ R ( F \otimes_R S) \leq \dim R/I .    
\end{equation*}
Finally, \eqref{Auslander_Buchsbaum} gives
\begin{equation*}
n\geq \projdim_R F = \depth_R S - \depth_R(F \otimes_R S) \geq \dim R - \dim R/I. 
\end{equation*}
Now, consider the case $s\geq 1$. We have the following sequence of (in)equalities
\begin{align} 
n &\geq \projdim_{R_\fp} F_\fp \notag\\
&=\depth_{R_\fp} S_\fp -\depth_{R_\fp}( F\otimes_R S)_\fp \notag\\
&= \depth_{R_\fp} S_\fp +s \label{inequality_with_p}\\
& \geq \depth_R S -\dim R/\fp +s \notag \\
&=\dim R- \dim R/\fp  +s. \notag
\end{align}
The first inequality is trivial, while the second follows from \eqref{depth_prime_ineq} applied to $S$, which is a derived $\fm$-complete $R$-complex. The first equality is from \eqref{Auslander_Buchsbaum}. The second equality follows from \eqref{depth_sup_ineq}, and the last one comes from $S$ being a big Cohen-Macaulay algebra over $R$. 

The proof of Theorem 2.2 in \cite{CF:2022} shows that $I\subseteq \fp$. Here are the details: we assume towards a contradiction that $I \nsubseteq \fp$. It follows that $F_\fp$ is isomorphic to $H_0(F)_\fp$ in the derived category, since $H_i(F)$ is $I$-power torsion for $i \geq 1$. This implies that $\sup H_*( F_\fp) = 0$, since we additionally have that $H(F_\fp)\neq 0$. We then have the following chain of (in)equalities 
\begin{align*}
\depth_{ R_\fp} R_\fp & = \depth_{R_\fp} F_\fp + \projdim_{R_\fp} F_\fp \\
&\geq \projdim_{R_p} F_\fp\\
&\geq \dim R - \dim R/\fp + s\\
&\geq \dim R_\fp + s. 
\end{align*} 
The equality is from \eqref{Auslander_Buchsbaum}. The first inequality is trivial, the second follows from \eqref{inequality_with_p}, and the last inequality is standard. Hence, we obtain $\depth R_\fp\geq \dim R_{\fp} +s$, which is a contradiction, since $s$ is positive. 

Therefore, we conclude that $I\subseteq \fp$, and from \eqref{inequality_with_p}, we obtain that
\begin{equation*}
n\geq \dim R -\dim R/I+s.    
\end{equation*}

Next, set $\Omega \coloneqq H_0(\Omega_{n-1}^S (F\otimes_R S))$. We claim that the $S$-module $\Omega$ is not projective. Indeed, since $s\leq n-1$, the $S$-complex 
\begin{equation*}
0 \to F_n \otimes_R S \to F_{n-1} \otimes_R S \to 0\,,    
\end{equation*}
with $F_{n-1} \otimes_R S$ in degree zero, is a free $S$-resolution of $\Omega$. Since $S$ is a big Cohen-Macaulay algebra, we have $\fm S \neq S$, where $\fm$ is the maximal ideal of $R$. Therefore, there exists a maximal ideal $\fn$ of $S$ containing $\fm S$. By $F_n\neq 0$, we have that
\begin{equation*}
\Tor[1] {S} {S/\fn} {\Omega}\cong (S/\fn)\otimes_S (F_n \otimes_R S) \cong (S/\fn)\otimes_R F_n \neq 0.    
\end{equation*}
Thus, $\Omega$ is not flat, and therefore not projective. We can now use \eqref{level_bound_ineq} to deduce that
\begin{equation*}
\level^S(F\otimes_R S)\geq n-s+1.
\end{equation*} 
From the base change result \cref{level_change_basis}, we then have 
\begin{equation*}
\level^R F \geq \level^S(F\otimes_R S)\geq n-s+1.    
\end{equation*}
Furthermore, from \eqref{inequality_with_I}, one has $n-s\geq \dim R-\dim R/I$, and the proof is complete. 
\end{proof}

Here is an immediate application of Theorem \ref{theorem_level_ineq}.

\begin{proposition}
Let $R$ be a commutative noetherian local ring.
\begin{enumerate}
\item \label{level_koszulcomplex_ideal} If $\underline{x}\coloneqq x_1, \ldots, x_n$ is a generating set for a proper ideal $I$ of $R$, then 
\begin{equation*}
\edim R+1 \geq \level^R K(\underline{x}; R)\geq \dim R -\dim R/I+1.    
\end{equation*}
\item \label{level_koszulcomplex_psop} If $\underline{x}= x_1, \ldots, x_n$ forms a  (partial) system of parameters for $R$, then
\begin{equation*}
\level^R K(\underline{x}; R)= \dim R -\dim R/I+1=n+1.    
\end{equation*}
\end{enumerate}    
\end{proposition}

\begin{proof}
For part \ref{level_koszulcomplex_ideal}, the inequality on the right comes from applying Theorem \ref{theorem_level_ineq} to the Koszul complex $K(\underline{x}; R)$ which is a perfect $R$-complex with $I=(\underline{x})$-torsion homology. For the inequality on the left, we take a minimal Cohen presentation of $R$, i.e., a surjective map $Q\twoheadrightarrow \hat{R}$, with $Q$ a regular ring and $\edim R=\dim Q$. From \cite[Theorem 5.5]{ABIM:2010} and $Q$ having finite global dimension, we deduce that
\begin{equation*}
\dim Q+1\geq \level^Q K(\underline{x}; Q).    
\end{equation*}
The exact functor $-\otimes^L_Q \hat{R}$ along with \eqref{level_change_basis} yield the following inequality
\begin{equation*}
\level^Q K(\underline{x}; Q)\geq \level^{\hat{R}} K(\underline{x}; \hat{R}).    
\end{equation*}
From \cite[Corollary 2.11]{L:2021} and the fact that the completion map $R\to \hat{R}$ is faithfully flat, we deduce that 
\begin{equation*}
\level^{\hat{R}} K(\underline{x}; \hat{R})=\level^{R} K(\underline{x}; R),    
\end{equation*}
which completes this part.

For part \ref{level_koszulcomplex_psop}, since $\underline{x}$ is a (partial) system of parameters, we have 
\begin{equation*}
\dim R-\dim R/I+1=n+1    
\end{equation*}
and it is enough to show that $\level^R K(\underline{x}; R)\leq \dim R-\dim R/I+1$. Note that the $R$-level of a finite free $R$-complex is always at most its length from \cite[Lemma 2.5.2]{ABIM:2010}. This yields the desired inequality
\begin{equation*}
\level^R K(\underline{x}; R)\leq n+1\,,    
\end{equation*}
which completes the proof. 
\end{proof}

\begin{remark} Part \ref{level_koszulcomplex_psop} of the previous proposition also demonstrates that the lower bound on the $R$-level provided by Theorem \ref{theorem_level_ineq} is optimal.  
\end{remark}

\subsection{Free rank} The \emph{free rank} of an $R$-module $M$ is the largest rank of a free direct summand of $M$; it is denoted by $\f_rank_{R}(M)$. We obtain the following lower bound on the level of the Koszul complex, and more generally over the setting of dg-algebras; consider \cite{A:2010} for the definition and properties of dg-algebras.

\begin{proposition} \label{level_frank_dgalgebra}
Let $(R,\fm, k)$ be a commutative noetherian local ring, $I$ a proper ideal in $R$, and $A$ a dg-algebra over $R$. If $H_0(A)=R/I$, and $A$ is a minimal dg-algebra, meaning $\partial_i(A)\subseteq \fm A_{i-1}$ for all $i$, then 
\begin{equation*}
\level^R A \geq \f_rank_{R/I} (I/I^2)+1.    
\end{equation*}
Specifically, when $I$ is generated by $(\underline{x})$, then 
\begin{equation*}
\level^R K(\underline{x};R) \geq \f_rank_{R/I} (I/I^2)+1.    
\end{equation*}
\end{proposition}

\begin{proof} 
The natural map $A\to H_0(A)=R/I$ can be extended to a morphism of complexes $f\colon A\to G$, where $G$ is a minimal free resolution of $R/I$ over $R$, with differentials denoted by $\partial^G$. Using \cite[Proposition 2.4]{AGMSV:2017}, it is enough to show that $\mathrm{Im} (f_s)\nsubseteq \fm G_s +\mathrm{Ker}(\partial_s^G)$, where $s=\f_rank_{R/I} (I/I^2)$. Since $G$ is a minimal free resolution, we have 
\begin{equation*}
\mathrm{Ker}(\partial_s^G)=\mathrm{Im}(\partial_{s+1}^G)\subseteq \fm G_s.    
\end{equation*}
Hence, $\mathrm{Im} (f_s)\nsubseteq \fm G_s +\mathrm{Ker}(\partial_s^G)$ is equivalent to $\mathrm{Im} (f_s)\nsubseteq \fm G_s$, which is in turn equivalent to $f_s \otimes k\neq 0$. We observe that for $\underline{x}\coloneqq x_1, \ldots, x_n$ a generating set of $I$, the induced map $K(\underline{x};R) \to R/I$ extends to a morphism of complexes $g\colon K(\underline{x};R)\to  G$. Also, specifying the image of each $x_i \in K_1(\underline{x}; R)$ in $A_1$ determines a map of dg $R$-algebras $h \colon K(\underline{x}; R) \to A$ and yields the following commutative diagram
\begin{equation*}
\begin{tikzcd}
K(\underline{x};R) \ar[rr, "h"] \ar[dr, "g" swap] & & A\ar[dl, "f"]\\
& G &
\end{tikzcd}    
\end{equation*}
By tensoring this diagram with the residue field $k$ and taking the $s$-th homology, we obtain the following commutative diagram
of dg-algebras 
\begin{equation*}
\begin{tikzcd}
H_s(\underline{x}; k) \ar[rr, "H_s(h\otimes k)"] \ar[dr, "H_s(g\otimes k)" swap] & & H_s(A\otimes_R^L k) \ar[dl, "H_s(f\otimes k)"]\\
& \Tor [s] R {R/I} k&
\end{tikzcd}    
\end{equation*}
It is easy to see that $H_1(g\otimes k)$ is an isomorphism. Therefore, the morphism $H_s(g\otimes k)$ factors as follows
\begin{equation*}
\begin{tikzcd}
H_s(\underline{x}; k) \ar[r, "\cong"] \ar[dr, "H_s(g\otimes k)" swap] &  \bigwedge^s \Tor[1] R {R/I} k \ar[d, "\kappa_s"]\\
& \Tor [s] R {R/I} k
\end{tikzcd}    
\end{equation*}
where $\kappa \colon \wedge \Tor[1] R {R/I} k\to\Tor R {R/I} k$ is the natural map of graded $k$-algebras. We claim that $\kappa_s\neq 0$, which implies that $H_s(f\otimes k)\neq 0$, and hence that $f_s \otimes k \neq 0$. 

To prove the claim, we first apply \cite[Proposition 2.1]{I:2001} to 
\begin{equation*}
\Tor R {R/I} k\cong R/I \otimes^L_R k   
\end{equation*}
and we get the following isomorphism of $k$-algebras
\begin{equation*}
\Tor R {R/I} k \cong B \otimes_{k} \Lambda    
\end{equation*}
where $B$ is a graded $k$-algebra with $B_0=k$ and
\begin{equation*}
\Lambda =\bigwedge (y_1, \ldots, y_s)\,,    
\end{equation*}
with $s=\f_rank_{R/I} (I/I^2)$ and $\partial y_i=0$, $|y_i|=1$ for all $i=1, \ldots ,s$. Then, $\Tor[1] R {R/I} k$ can be realized as $B_1 \oplus (y_1, \ldots, y_s)$, and by taking exterior algebras on the map $(y_1, \ldots, y_s)\hookrightarrow \Tor[1] R {R/I} k$ the following embedding is being deduced
\begin{equation*}
\Lambda=\bigwedge (y_1, \ldots, y_s) \hookrightarrow \bigwedge \Tor[1] R {R/I} k.  
\end{equation*}
Next, by composing with $\kappa$, we get the map
\begin{equation*}
\Lambda \hookrightarrow \bigwedge \Tor[1] R {R/I} k \xrightarrow{\kappa} \Tor R {R/I} k= B \otimes_k \Lambda\,,    
\end{equation*}
which is clearly non-zero on degree $s$, as $\Lambda_s=\bigwedge^s (y_1, \ldots, y_s) \neq 0$. Hence, $\kappa_s\neq 0$, and the proof has been completed.
\end{proof}

\subsection{Lech-Independent sequences} A sequence $\underline{x}=x_1, \ldots, x_n$ is called \emph{Lech-independent} if for $I=(\underline{x})=(x_1,\ldots, x_n)$, the natural surjection
\begin{equation*}
(R/I)^n \longrightarrow I/I^2    
\end{equation*}
is an isomorphism, see \cite{B:2023, Ha:2001, L:1964}. A regular sequence is also Lech-independent, and the converse holds when, additionally, $I$ has finite projective dimension, see \cite{V:1967}. Another example of a Lech-independent sequence is a minimal generating set of the maximal ideal of a local ring.  Lech-independent sequences are closed under flat base change.

The following result is immediate from \cref{level_frank_dgalgebra} and generalizes the \cite[Theorem 4.2(1)]{AGMSV:2017}, which concerns the case of the maximal ideal. 

\begin{corollary} \label{level_koszulcomplex_indep} Let $(R,\fm, k)$ be a commutative noetherian local ring and $\underline{x}=x_1,\ldots, x_n$ be a Lech-independent sequence, then 
\begin{equation*}
\level^R K(\underline{x};R) =n+1.    
\end{equation*}
\end{corollary}

\begin{proof}
It is clear that $\level^R K(\underline{x}; R)\leq n+1$ from \cite[Lemma 2.5.2]{ABIM:2010}. For the reverse inequality, we apply \cref{level_frank_dgalgebra} to the sequence $\underline{x}=x_1, \ldots, x_n$. From $\underline{x}$ being a Lech-independent sequence, $\f_rank_{R/I}(I/I^2)=n$, and this completes the proof. 
\end{proof}

\section{Examples}
With stronger assumptions on the complex, a sharper lower bound than the one in \cite[Theorem 4.2]{AIN:2018}—involving the superheight of an ideal—was established in \cite[Theorem 5.1]{ABIM:2010}. The \emph{superheight} of an  ideal $I$ is defined as the number 
\begin{equation*}
\superheight I = \sup \set{\height IT}{ T \text{ is a noetherian } R\text{-algebra}}.   
\end{equation*}
The following result is a special case of \cite[Theorem 5.1]{ABIM:2010} and has also been proved in \cite[Theorem 3.2]{AGMSV:2017}. Using Theorem \ref{theorem_level_ineq}, the proof is significantly simplified. 

\begin{corollary}
Let $R$ be a commutative noetherian local ring,  $F$ a perfect $R$-complex and $I\subseteq R$ the annihilator of $\bigoplus_{i\in \mathbb{Z}} H_i(F)$. Then, the following inequality holds
\begin{equation*}
\level^R F \geq \superheight I + 1.
\end{equation*}
\end{corollary}

\begin{proof}
Let $S$ be a noetherian $R$-algebra. The ideal $IS$ annihilates 
\begin{equation*}
\bigoplus_{i \in \mathbb{Z}} H_i(S \otimes F)\,,    
\end{equation*}
hence, by applying the previous theorem, we get the first inequality
\begin{equation*}
\level^S(S\otimes F) \geq \dim S -\dim S/IS +1\geq \height I+1,
\end{equation*}
the second one is standard. Additionally, from \eqref{level_change_basis}, we obtain that 
\begin{equation*}
\level^R F\geq  \level^S(S\otimes F)\,,    
\end{equation*}
which completes the proof.
\end{proof}

Another interesting number is the \emph{bigheight} of an ideal $I$ defined as follows:
\begin{equation*}
\bigheight I = \sup \set{\height \fp }{\fp \text{ is minimal prime over } I}.    
\end{equation*}
It is easy to check that we always have the following inequalities
\begin{equation*}
\height I\leq \bigheight I\leq \superheight I
\end{equation*}
and that the first one can be strict. Additionally, it was shown in \cite{K:1988} that the second inequality can also be strict. We recall that the following inequality always holds 
\begin{equation*}
\height I\leq \dim R-\dim R/I.    
\end{equation*}
However, we observe that the quantities $\dim R-\dim R/I$ and $\bigheight I$ (or even $\superheight I$) cannot be compared. 

On the one hand, take the ring $R=k\llbracket x_1, \ldots, x_n\rrbracket$ and the ideal $I=(x_1)\cap(x_2,\ldots, x_n)$ for which we have 
\begin{equation*}
1=\dim R-\dim R/I \quad \text{and}\quad \bigheight I=n-1.    
\end{equation*}
On the other hand, the ring $R=k\llbracket x_1, \ldots, x_n \rrbracket/ ((x_1)\cap (x_2,\ldots,x_n))$ and the ideal $I=(x_2, \ldots, x_n)$ give
\begin{equation*}
n-2=\dim R-\dim R/I \quad \text{and}\quad \bigheight I=0.    
\end{equation*}
This last example also shows that the quantity $\dim R-\dim R/I$ can be arbitrarily larger than $\height I$. 

\begin{remark}
One could ask whether, under the conditions of Theorem \ref{theorem_level_ineq}, the quantity $\dim R - \dim R/I$ could be replaced by $\superheight I $ or $\bigheight I$. The answer is negative, as we can see by considering the ring $R=k\llbracket x_1, \ldots, x_n \rrbracket$ for $n\geq 3$, the ideal $I=(x_1)\cap (x_2,\ldots, x_n)$ and the perfect $R$-complex $F=K(x_1; R)$. It is clear that $F$ has $I$-torsion homology and $\level^R F=2$, but $\bigheight I=n-1$.
\end{remark}

\subsection{Tensor nilpotent and fiberwise zero maps} A morphism $f\colon X \to Y$ in $\cat{D}(R)$ is called \emph{tensor nilpotent} if for some $n \in \mathbb{N}$ the map $\otimes^n f\colon \otimes^n_R X \to \otimes^n_R Y$ is zero in $\cat{D}(R)$. A morphism $f\colon X \to Y$ in $\cat{D}(R)$ is called \emph{fiberwise zero} if for all $\fp \in \Spec(R)$ the map $k(\fp)\otimes^L_R f$ is zero in $\cat{D}(k(\fp))$; see \cite[3.2]{AIN:2018}. These two notions are equivalent when the map $f$ is between perfect $R$-complexes, see \cite{H:1987,N:1992}. In this subsection, we will investigate the relation between Theorem \ref{theorem_level_ineq} and \cite[Theorem 4.1]{AIN:2018}. This result states that over a commutative noetherian ring $R$, if there exists a morphism $f \colon G \to F$ of perfect $R$-complexes, which is not fiberwise zero and factors through an $R$-complex with $I$-torsion homology for some ideal $I \subseteq R$, then the following inequality holds
\begin{equation}
\label{theorem_AIN_4.1}
\level^R \Hom R G F \geq  \height I + 1.    
\end{equation}
This result led to the version \cite[Theorem 4.2]{AIN:2018} of the Improved New Intersection Theorem. We observe that under the conditions of \ref{theorem_AIN_4.1}, we cannot replace $\height I$ with $\dim R-\dim R/I$. To see this, we take the ring 
\begin{equation*}
R=k \llbracket x_1, \ldots, x_n \rrbracket/((x_1)\cap (x_2,\ldots,x_n))\,,    
\end{equation*}
the perfect $R$-complexes $F=G=R$ and the ideal $I=(x_2, \ldots, x_n)$. Consider the map $f\colon R\xrightarrow{x_1} R$, which is not tensor nilpotent since $x_1$ is not a nilpotent element in $R$. From the following diagram, one can observe that the map $f$ can be factored through the $I$-torsion $R$-complex $K(x_2, \ldots, x_n; R)$.
\begin{equation*}
\begin{tikzcd}
& & & &  0 \ar[r] \ar[d] & R \ar[r] \ar[d,"\mathrm{id}"] & 0 \\
0 \ar[r] & R \ar[r] & R^{n-1} \ar[r]  & \cdots \ar[r] & R^{n-1} \ar[r]  \ar[d] & R \ar[r] \ar[d,"x_1"] & 0 \\
& & & & 0 \ar[r] & R \ar[r]  & 0
\end{tikzcd}
\end{equation*}    
We see that  $\level^R \Hom R R R=\level^R R=1$, while the quantity 
\begin{equation*}
\dim R-\dim R/I +1= n-1    
\end{equation*}
can be arbitrarily large. 

Another natural question is whether, in \ref{theorem_AIN_4.1}, one could replace $\height I$ with $\bigheight I$. The answer is negative, as we can see by considering the ring $R=k \llbracket x_1, \ldots, x_n \rrbracket$, the ideal $I=(x_1)\cap (x_2,\ldots, x_n)$, and the perfect $R$-complexes $F=G=K(x_1; R)$. Then, we observe that the identity map on $F$ factors through the $I$-power torsion $R$-complex, $R\Gamma_I F$, and is not a fiberwise zero map. In this case, we have 
\begin{equation*}
\level^R \Hom R F F =\level^R F=2,    
\end{equation*}
but $\bigheight I +1 =n$. 

\bibliographystyle{amsplain}

\end{document}